OpenStudy (anonymous):

Which statement best describes the graph of -x3 - x2 + 4x + 4? It starts down on the left and goes up on the right and intersects the axis at x = -2, 2, and 4. It starts down on the left and goes up on the right and intersects the axis at x = -2, -1, and 2. It starts up on the left and goes down on the right and intersects the x-axis at x = -2, -1, and 2. It starts up on the left and goes down on the right and intersects the axis at x = -2, 2, and 4.

OpenStudy (anonymous):

please help me...

OpenStudy (anonymous):

\[\begin{align*}-x^3-x^2+4x+4&=-x^2(x+1)+4(x+1)\\&=(4-x)^2(x+1)\end{align*}\] So you know the zeroes of the polynomial are \(x=-1,4\).  To figure out the behavior of the polynomial around these zeroes, pick a value from the spaces between the zeroes and plug them in.